OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3

OpenStudy (texaschic101):

do you know how to do the substitution method ?

OpenStudy (anonymous):

no i dont will u please help me?

OpenStudy (caozeyuan):

y=x-3, so 4x+2(x-3)=12, 6x-6=12,6x=18,x=3. Solved!

OpenStudy (anonymous):

thank you that was alot of help

OpenStudy (texaschic101):

That is correct